# Lifted Forward Planning in Relational Factored Markov Decision Processes with Concurrent Actions

### v2

@@ 3.1. Parameterizing Actions
In our epidemic example, the mayor, representing the government, can impose travel bans on all parts of the town’s population. We extend the action definition in this subsection to account for groups of objects. Having groups, we circumvent enumeration of all possible subsets and model the example action of imposing travel bans on a subset of the population efficiently.
@@
Action PRVs allow for a more general action setting. When writing action, we refer to Definition 3.1, i.e., concurrent actions. In our example, the mayor can restrict multiple persons from travelling at once:
@@
When the persons $a$ to $h$ are indistinguishable, it is irrelevant which permutation is used for the constraints, only the amounts are relevant, which we exploit in Foreplan.
As actions are now parameterized, we describe the impact of parameterization on the rewards.
@@ 3.2. Parameterized Local Reward Functions
The reward function in fMDPs maps from the (joint) state to the reward of the state. For evaluating the reward function, we thus have to construct the joint state and cannot exploit the factorization, breaking our aim of compact representation. To further use our compact representation, we introduce a decomposable reward function: For simplicity, we assume that the reward function is factored as $R=\sum_{i}R_{i}$, with local reward functions $R_{i}$ with scope restricted to a subset of the state variables. Other operations are possible too, as long as they have a hyperoperation, as already mentioned. As we have indistinguishable state variables, we reduce redundant computations by using representatives in the reward functions analogue to a parfactor, but using a sum instead of a product:
@@ 4. Foreplan: Efficient Forward Planning
In this section, we propose Foreplan, our exact forward relational planner for rfMDPs. The input to Foreplan is an rfMDP. The output is a value function, which induces a policy. Foreplan computes the value function by using a compact state representation for the rfMDP to exploit the indistinguishability of the objects and then running a linear program based on the state representation to calculate the value function.
The first-order representation in rfMDP contains dependencies between PRVs. For using the representation in computation, we first need to recognize the dependencies and find a suitable representation of them to compute the value function. We first describe how Foreplan recognizes the dependencies and identifies a suitable representation, Afterwards, we outline how Foreplan uses the identified representation to compute a value function.
@@ 4.1. Obtaining a Compact State Representation
Foreplan needs to encode the current state compactly to efficiently reason about indistinguishable variables. Thus, in this subsection, we describe how Foreplan treats indistinguishable objects for computations using rfMDPs. The representation in rfMDPs does not consider dependencies between PRVs. But, Foreplan can tame the information to store: Foreplan does not need to keep track of objects that could be differentiated by their history. Rather, with each action and new time step, the history is swept away and the objects remain indistinguishable because of the first-order Markov assumption. Thus, the basic idea is to count the objects for which the PRVs are true or false, which is inspired by Counting Random Variables (CRVs) crv. When evaluating a parfactor, the joint assignment to the PRVs is needed, rather than individual counts for each PRV. It is sufficient to count the number of occurrences of each possible truth-value assignment to the groundings of the input PRVs of a parfactor in a histogram. The input PRVs of a parfactor are the ones representing the current state. Counting only the input PRVs is sufficient, because all possible next states are iterated separately, which we describe in the next subsection. But some parfactors may be defined over PRVs defined over different logvars, like $f_{3}$ in Figure 2 is defined over $Travel(M)$ and $Epidemic$. Given the counts for $Travel(M)$ and $Epidemic$ separately is enough, since these PRVs do not depend on each other. But how does Foreplan recognize which PRVs depend on which PRVs? For answering this question, we propose the relational cost graph to identify the dependencies between PRVs and identify the groups that need to be counted jointly. Focusing only on the counts for the groups enables Foreplan to use a much simpler state space representation, namely the set of possible histograms, compared to the grounded state space representation, which is the Cartesian product over all state variables. We now describe in more detail how to count the assignments.
Counting the assignments of each PRV separately is insufficient, as PRVs can be defined over the same logvars and thus interfere with each other. However, the parfactors can be evaluated separately since, in Equation 2, we have the full current and next state available. Thus, it is sufficient to count PRVs together if they share a logvar and occur in a parfactor together. To obtain the representation and later on quantify its complexity, we define the relational cost graph:
@@
The relational cost graph for Example 3 consists of three isolated vertices corresponding to $Sick(M)$, $Travel(M)$ and $Epidemic$. The first two do not occur together in a parfactor or local reward function and the last does not share a logvar with one of the first two. Figure 3 shows the visual representation of the relational cost graph.
@@
The key insight now is that (maximal) cliques in the relational cost graph correspond to sets of PRVs that Foreplan needs to count together as they interfere with each other. But when having more than one logvar in a PRV, basic CRVs crv only allow to count one of them. To use our representation in computations, we extend CRVs to allow for more than one logvar, extending the definition by crv:
@@
A counting formula $\gamma=\#_{C}[B_{1},\dots,B_{k}]$ is defined over PRVs $B_{i}$ with a constraint $C=(\mathcal{L},C_{\mathcal{L}})$ over the logvars $\mathcal{L}$ of the PRVs $B_{i}$. The counting formula represents a counting random variable (CRV) whose range is the set of possible histograms that distribute $n=\left\lvert C_{\mathcal{L}}\right\rvert$ elements into $\prod_{i=1}^{k}\left\lvert range(B_{i})\right\rvert$ buckets. The state of $\gamma$ is the histogram function $h=\{(r_{i},n_{i})\}_{i=1}^{r}$ stating for each $r_{i}\in\times_{i=1}^{k}range(P_{i})$ the number $n_{i}$ of tuples $(P_{i})_{i}$ whose state is $r_{i}$.
If no restrictions apply, we omit $C$, and $n=\prod_{L\in\mathbf{L}}\left\lvert\mathcal{D}(L)\right\rvert$, where $\mathbf{L}$ is the set of logvars of the PRVs $B_{i}$. We do not define the operations to manipulate CRVs as we do not need them in this paper. The CRV corresponding to a clique gives us the number of occurrences for each possible instantiation of the PRVs in that clique. We give a small example for a CRV using an additional PRV $Friends(M,Y)$, with $M$ and $Y$ having the same domain, for illustrative purposes:
@@
To have a PRV with two logvars, assume we have three PRVs, $Travel(M)$, $Friends(M,Y)$ and $Sick(Y)$ with $\left\lvert\mathcal{D(M)}\right\rvert=\left\lvert\mathcal{D}(Y)\right\rvert=8$. A possible state for the CRV #[Travel(M)\#[Travel(M), $Friends(M$, $Y)$, $Sick(Y)]$ is $\{(ttt,6)$, $(ttf,30)$, $(tft,4)$, $(tff,0)$, $(ftt,6)$, $(ftf,0)$, $(fft,0)$, (fff,18)}(fff,18)\} and we have $n=64$. The first histogram tells, that in this state, there are six tuples for which $Travel(a)=t$, $Friends(a,b)=t$, and $Sick(b)=t$ holds.
Example 4.1 illustrates that we need only eight buckets in a histogram regardless of the domain size of $M$ and $Y$. The number of buckets depends on the range values of the PRVs and not on the domain sizes of them. We now formalize the state representation, which is basically a set of CRVs per (maximal) clique in the relational cost graph:
@@
We create one CRV $S_{i}$ for each (maximal) clique in the relational cost graph, counting the corresponding PRVs in the clique. For a single propositional RV in the relational cost graph, we use the RV for $S_{i}$ instead of a CRV. A state assigns a value to each $S_{i}$. The resulting state space is the set of possible states. We denote the state space by the tuple $(S_{i})_{i}$, which contains one CRV (or RV) per clique.
@@
The relational cost graph in Example 4.1 tells us that we can count the number of healthy persons separately from the number of travelling persons. In particular, we do not store which sick persons are travelling, since we do not need this information. We prove that our state representation is correct, i.e., retains the same semantics, namely exactly representing $\mathbf{S}$: {theorem} The representation in Definition 4.1 is correct.
@@
Given groundings for the state PRVs, we derive the histograms for the CRVs by counting the assignments for each parfactor. Given a representation as defined in Definition 4.1, we reconstruct, for each parfactor, the groundings of the PRVs by extracting the counts from the CRVs and instantiating the respective parfactors. We provide a full proof in Appendix A.4. ∎
@@
The action $Restrict(M)$ uses the PRV $Travel(M)$. Thus, the mayor needs to specify how many persons of those (not) travelling are (not) allowed to travel. Suppose that in the current state five out of eight persons are travelling and two out of eight persons are sick. The action $Restrict(M)$ is defined over #[Travel(M),Restrict(M)]\#[Travel(M),Restrict(M)]. A concrete action is, e.g., $a=\{(tt,3)$, (ft,2)}(ft,2)\}, which means that three persons currently travelling are restricted from travelling and two persons not travelling. The action $a$ does not need to specify the counts of people no travel ban is imposed on ($tf$, ffff), as these are determined by $a$ and the current state.
@@
4.2. Computing a Value Function
@@
For instantiating the linear program in Equation 2, it is left to describe how Foreplan calculates the transition probability. Since we have full evidence provided, Foreplan evaluates each state CRV $S_{i}^{\prime}$ separately. For a fixed state CRV $S_{i}^{\prime}$, the value $s_{i}^{\prime}$ is fixed since the whole state space is iterated. For evaluating $P(s_{i}^{\prime}\mid s,a)$, Foreplan iterates over all possible assignment transitions, e.g., from each bucket in the histogram $s$ to each bucket in the histogram $s^{\prime}$. In our epidemic example, this is, e.g., the number of people getting sick and healthy. For each possible assignment transition, Foreplan calculates the transition probability: The probability for a representative transition is given by the involved parfactors, i.e., the parfactors defined over the PRVs mentioned in $S_{i}^{\prime}$. But this representative probability has to be weighted by how many times this assignment transition is applicable using the multinomial coefficient. Take for example the state in which five persons are sick. When three sick persons get healthy, this assignment transition has weight $5\choose 3$ as any three of the five persons can get healthy. The final probability of $P(s_{i}^{\prime}\mid s,a)=\sum_{t}w_{t}p_{t}$ for a fixed CRV is the sum over all possible assignment transitions $t$, each with representative probability $p_{t}$ and weight $w_{t}$. Since the state space is factored, the full transition probability $P(s^{\prime}\mid s,a)$ is factored as $P(s^{\prime}\mid s,a)=\prod_{i}P(s_{i}^{\prime}\mid s_{i},a)$. We provide a more detailed description and an example in Appendix A.5.
@@ 5. Complexity Analysis of Foreplan
For each clique, the size of the histogram function is exponential in the number of vertices in the clique, as we enumerate all possible assignments. Thus, the size of the state representation is bounded by $c\times 2^{w}$. ∎
Note that $c$ and $w$ are independent of the domain sizes and determined only by the structure of the relational cost graph, and thus in general small. Also, $c$ is bounded by the number of PRVs in our parfactor model, as one parfactor is sufficient per PRV. Theorem 5 overapproximates the size of the state representation, as not all cliques have the same size and not both, $c$ and $w$ are large at the same time. Building on the size of the state representation, we give the complexity of the state and action space: {theorem} The state and action spaces are both polynomial in the number of objects and exponential in $c$ and $w$.
@@
The number of linear programming constraints and variables are polynomial in the state space.
@@
By Equation 2, Foreplan generates one variable per possible state and one constraint for each state and action combination. ∎
Plugging Theorem 5 in Theorem 5 leads to: {theorem} The runtime of Foreplan is polynomial in the number of objects and exponential in $c$ and $w$.
Therefore, with Foreplan, we have introduced an exact planner with a runtime polynomial instead of exponential w.r.t domain sizes for concurrent actions. The other parameters Foreplan is exponential in are mostly small and fixed. Thus, we have already achieved an exponential speed up.
Our main contribution Foreplan greatly advances the state of the art. However, as it is an exact algorithm, the value function is still not factorized. In case the runtime is of utter importance, we can also approxiamte the value function, following the idea of approx_factored_mdps, which has also been picked up by, e.g., 10.5555/3020336.3020398. By approximating the value function, we can prevent iterating over the joint state space leading to a blazingly fast, but approximate, version of Foreplan.
@@ 6. Foreplan: Even Faster by Approximation
While Foreplan runs in time polynomial in the number of objects, the runtime still depends exponentially on $c$. In this section, we present an approximation technique inspired by the Approximate Linear Programming (ALP) approx_factored_mdps approach to prevent iterating the whole state space. Our approach follows the same idea as ALP and ALP for first-order MDPs 10.5555/3020336.3020398. We first describe the approximation idea and then how Foreplan uses the approximation for our new representation and for concurrent actions. Last, we give bounds on the runtime and on the approximation quality. We call the approximate version Approximate Foreplan. Since Approximate Foreplan is an extension of our main contribution Foreplan, we refer to the Appendix for technical details.
Foreplan needs to iterate over the whole state space, because the value function maps from a state to the value of that state. We approximate the value function by a set $h_{i}$ of basis functions, whose scope is a subset of $\mathbf{S}$, $V\approx\sum_{i}w_{i}h_{i}$, where the goal is to find the most suitable weights $w_{i}$ approx_factored_mdps. Approximate Foreplan also needs the value of all possible next states in terms of the same approximation. Therefore, Approximate Foreplan makes use of the backprojections $g_{i}^{a}$ of the basis functions $h_{i}$ approx_factored_mdps, stating the influence of $x$ on the next state:
@@
To compute the basis functions and backprojections lifted, we parameterize the basis functions analogue to the reward functions and calculate them lifted in the same way. The basis functions should capture the important dynamics in the model koller1999computing, most importantly the rewards 10.5555/3020336.3020398. As proposed by koller1999computing and picked up in 10.5555/3020336.3020398, we also use basis functions for capturing the rewards alongside a constant basis function:
@@
The backprojections are computed lifted as defined next:
@@
Approximate Foreplan precomputes all backprojections and then builds the following linear program approx_factored_mdps, whose result is the approximated value function and constitutes the result of Approximate Foreplan:
@@
The $\alpha_{i}$’s are effectively coefficients for a linear combination over the $w_{i}$, stating how important the minimization of each $w_{i}$ is approx_factored_mdps; doi:10.1287/opre.51.6.850.24925. The maximum operator is not part of the definition of linear programs and is removed in an operation similar to variable elimination (VE) ve. In Appendix A.6, we provide an example for the removal procedure. Moreover, in Appendix B, we show a walkthrough of Approximate Foreplan on the epidemic example.
@@
Most notably, $w$ and induced width in Theorem 6 are mostly small and fixed, leading to a polynomial runtime, as the growth in the number of objects is of more interest. In Appendix A.8, we combine the relational cost graph and the cost networks in a single total relational cost graph, and show that the runtime Approximate Foreplan is polynomial in the number of objects when the treewidth of the total relational cost graph is bounded.
Moreover, we can show that Approximate Foreplan and ALP compute the same solutions: {theorem} Given an rfMDP $R$, Approximate Foreplan and ALP are equivalent on $R$ and the grounded version of $R$, respectively.
@@
The full proof is in Appendix A.9. The objective function groups grounded basis functions and calculates them lifted when choosing appropiate $\alpha_{i}$. Each individual constraint is correct, because the lifted backprojections and lifted basis functions compute the same value as for the grounded model. The action representation covers the whole action space. ∎
With ALP and Approximate Foreplan being equivalent, we can transfer the approximation guarantee for ALP to Approximate Foreplan:

### v3

@@ 3.1. Parameterizing Actions
In our epidemic example, the mayor, representing the government, can impose travel bans on all parts of the town’s population. We extend the action definition in this subsection to account for groups of objects. Having groups, we circumvent enumerating of all possible subsets and model the example action of imposing travel bans on a subset of the population efficiently.
@@
Action PRVs allow for a more general action setting. When writing action, we refer to Definition 3.1. In our example, the mayor can restrict multiple persons from travelling at once:
@@
However, when the persons $a$ to $h$ are indistinguishable, it is irrelevant which selection is used for the constraints, only the amounts are relevant. We describe the impact of parameterization on the rewards next.
@@ 3.2. Parameterized Local Reward Functions
The reward function in fMDPs maps from the (joint) state to the reward of the state. For evaluating the reward function, we thus have to construct the joint state and cannot exploit the factorization, breaking our aim of compact representation. To further use our compact representation, we introduce a decomposable reward function: For simplicity, we assume that the reward function is factored as $R=\sum_{i}R_{i}$, with local reward functions $R_{i}$ with scope restricted to a subset of the state variables. Other operations are possible too, as long as they have a hyperoperation, as already mentioned. As we have indistinguishable state variables, we reduce redundant computations by using representatives in the reward functions analogously to a parfactor, but using a sum instead of a product:
@@ 4. Foreplan: Efficient Forward Planning
In this section, we propose Foreplan, our exact forward relational planner for rfMDPs. The input to Foreplan is an rfMDP. The output is a value function, which induces a policy. Foreplan computes the value function using a compact state representation for the rfMDP to exploit the indistinguishability of the objects and then running a linear program based on the state representation to calculate the value function.
The first-order representation in rfMDP contains dependencies between PRVs. For using the representation in computation, we first need to recognize the dependencies and find a suitable representation of them to compute the value function efficiently. We first describe how Foreplan recognizes the dependencies and identifies a suitable representation. Afterwards, we outline how Foreplan uses the identified representation to compute a value function.
@@ 4.1. Obtaining a Compact State Representation
Foreplan needs to encode the current state compactly to efficiently reason about indistinguishable variables. Thus, in this subsection, we describe how Foreplan treats indistinguishable objects for computations using rfMDPs. Foreplan does not need to keep track of objects that could be differentiated by their history. Rather, with each action and new time step, the history is swept away and the objects remain indistinguishable because of the first-order Markov assumption. To compactly represent the state, the basic idea is to count the objects for which the PRVs are true or false, using the idea of Counting Random Variables (CRVs) crv. When evaluating a parfactor, the joint assignment to the PRVs is needed, rather than individual counts for each PRV. It is sufficient to count the number of occurrences of each possible truth-value assignment to the groundings of the input PRVs of a parfactor in a histogram. The input PRVs of a parfactor are the ones representing the current state. Counting only the input PRVs is sufficient, because all possible next states are iterated separately later. But some parfactors may be defined over PRVs defined over different logvars, like $f_{3}$ in Figure 2 is defined over $Travel(M)$ and $Epidemic$. Given the counts for $Travel(M)$ and $Epidemic$ separately is enough, since these PRVs do not depend on each other. But how does Foreplan recognize which PRVs depend on which PRVs? For answering this question, we propose the relational cost graph to identify the dependencies between PRVs and identify the groups that need to be counted jointly. Focusing only on the counts for the groups enables Foreplan to use a much simpler state space representation, namely the set of possible histograms, compared to the grounded state space representation, which is the Cartesian product over the domains of all state variables. We now describe how to count the assignments in more detail.
Counting the assignments of each PRV separately is insufficient, as PRVs can be defined over the same logvars and thus interfere with each other. However, the parfactors can be evaluated separately since, in Equation 2, we have the full current and next state available. Thus, it is sufficient to count PRVs together if they share a logvar and occur together in a parfactor. To obtain the representation and later on quantify its complexity, we define the relational cost graph:
@@
The relational cost graph for Example 3 consists of three isolated vertices corresponding to $Sick(M)$, $Travel(M)$ and $Epidemic$. The first two do not occur together in a parfactor or local reward function and both do not share a logvar with the last one. Figure 3 shows the visual representation of the relational cost graph.
@@
The key insight now is that (maximal) cliques in the relational cost graph correspond to sets of PRVs that Foreplan needs to count together as they interfere with each other. For one logvar per PRV, basic CRVs crv already return the result. To lift the limitation of one logvar per PRV, we extend the definition by crv:
@@
A counting formula $\gamma=\#_{C}[B_{1},\dots,B_{k}]$ is defined over PRVs $B_{i}$ with a constraint $C=(\mathcal{L},C_{\mathcal{L}})$ over the logvars $\mathcal{L}$ of the PRVs $B_{i}$. The counting formula represents a counting random variable (CRV) whose range is the set of possible histograms that distribute $n=\left\lvert C_{\mathcal{L}}\right\rvert$ elements into $\prod_{i=1}^{k}\left\lvert\mathcal{R}(B_{i})\right\rvert$ buckets. The state of $\gamma$ is the histogram function $h=\{(r_{i},n_{i})\}_{i=1}^{r}$ stating for each $r_{i}\in\times_{i=1}^{k}\mathcal{R}(B_{i})$ the number $n_{i}$ of tuples $(B_{i})_{i}$ whose state is $r_{i}$.
If no restrictions apply, we omit $C$, and $n=\prod_{L\in\mathbf{L}}\left\lvert\mathcal{D}(L)\right\rvert$, where $\mathbf{L}$ is the set of logvars of the PRVs $B_{i}$. We do not define the operations to manipulate CRVs as we do not need them in this paper. The CRV corresponding to a clique gives us the number of occurrences for each possible instantiation of the PRVs in that clique. For illustrative purposes, we give a small example for a CRV using an additional PRV $Friends(M,Y)$, with $M$ and $Y$ having the same domain:
@@
To have a PRV with two logvars, assume we have three PRVs, $Travel(M)$, $Friends(M,Y)$ and $Sick(Y)$ with $\left\lvert\mathcal{D(M)}\right\rvert=\left\lvert\mathcal{D}(Y)\right\rvert=8$. A possible state for the CRV #[Travel(M)\#[Travel(M), $Friends(M$, $Y)$, $Sick(Y)]$ is $\{(ttt,6)$, $(ttf,30)$, $(tft,4)$, $(tff,0)$, $(ftt,6)$, $(ftf,0)$, $(fft,0)$, (fff,18)}(fff,18)\} and we have $n=64$. The first histogram entry shows that, in this state, there are six tuples for which $Travel(m)=t$, $Friends(m,y)=t$, and $Sick(y)=t$ holds.
Example 4.1 illustrates that we need only eight buckets in a histogram regardless of the domain size of $M$ and $Y$. The number of buckets depends on the range values of the PRVs and not on the domain sizes. We formalize the state representation, which is a set of CRVs per (maximal) clique in the relational cost graph:
@@
For counting the corresponding PRVs in a clique, we create one CRV $S_{i}$ for each (maximal) clique in the relational cost graph. For a single propositional RV in the relational cost graph, we use the RV for $S_{i}$ instead of a CRV. A state assigns a value to each $S_{i}$. The resulting state space is the set of possible states. We denote the state space by the tuple $(S_{i})_{i}$, which contains one CRV (or RV) per clique.
@@
The relational cost graph in Example 4.1 tells us that we can count the number of healthy persons separately from the number of travelling persons. In particular, Foreplan does not store which sick persons are travelling, since this information is irrelevant. We prove that our state representation is correct, i.e., retains the same semantics, namely exactly representing $\mathbf{S}$ of a ground MDP: {theorem} The representation in Definition 4.1 is correct.
@@
Given groundings for the state PRVs, we derive the histograms for the CRVs by counting the assignments for each parfactor. Given a representation as in Definition 4.1, we reconstruct, for each parfactor, the groundings of the PRVs by using the counts from the CRVs to instantiate the respective parfactor. We provide a full proof in Appendix A.4. ∎
@@
The action $Restrict(M)$ uses the PRV $Travel(M)$. Thus, the mayor needs to specify how many persons of those (not) travelling are (not) allowed to travel. Suppose that in the current state five out of eight persons are travelling and two persons are sick. The action $Restrict(M)$ is defined over #[Travel(M),Restrict(M)]\#[Travel(M),Restrict(M)]. A concrete action is, e.g., $a=\{(tt,3)$, (ft,2)}(ft,2)\}, which means that three persons currently travelling are restricted from travelling and two persons not travelling. The action $a$ does not need to specify the counts of people no travel ban is imposed on ($tf$, ffff), as these are determined by $a$ and the current state.
@@
4.2. Computing the Value Function
@@
For instantiating the linear program in Equation 2, we next describe how Foreplan calculates the transition probability. Since we have full evidence provided, Foreplan evaluates each state CRV $S_{i}^{\prime}$ separately. For a fixed state CRV $S_{i}^{\prime}$, the value $s_{i}^{\prime}$ is fixed since the whole state space is iterated. For evaluating $P(s_{i}^{\prime}\mid s,a)$, Foreplan iterates over all possible assignment transitions, e.g., from each bucket in the histogram $s$ to each bucket in the histogram $s^{\prime}$. In our epidemic example, this is, e.g., the number of people getting sick and healthy. For each possible assignment transition, Foreplan calculates the transition probability: The probability for a representative transition is given by the involved parfactors, i.e., the parfactors defined over the PRVs mentioned in $S_{i}^{\prime}$. But this representative probability has to be weighted by how many times this assignment transition is applicable using the multinomial coefficient. Take for example the state in which five persons are sick. When three sick persons get healthy, this assignment transition has weight $5\choose 3$ as any three of the five persons can get healthy. The final probability of $P(s_{i}^{\prime}\mid s,a)=\sum_{t}w_{t}p_{t}$ for a fixed CRV is the sum over all possible assignment transitions $t$, each with representative probability $p_{t}$ and weight $w_{t}$. Since the state space is factored, the full transition probability $P(s^{\prime}\mid s,a)$ is factored as $P(s^{\prime}\mid s,a)=\prod_{i}P(s_{i}^{\prime}\mid s_{i},a)$. We provide a more detailed description and an example in Appendix A.5.
@@ 5. Complexity Analysis of Foreplan
For each clique in the relational cost graph, the size of the histogram function is exponential in the number of vertices in the clique, as we enumerate all possible assignments. Thus, the size of the state representation is bounded by $c\times 2^{w}$. ∎
Note that $c$ and $w$ are independent of the domain sizes and determined only by the structure of the relational cost graph, and thus in general small. Also $w$ is bounded by the number of PRVs and $c$ by the number of parfactors in the parfactor model. Theorem 5 overapproximates the size of the state representation, as not all cliques have the same size and not both, $c$ and $w$ are large at the same time. Building on the size of the state representation, we give the complexity of the state and action space: {theorem} The state and action spaces are both polynomial in the number of objects and exponential in $c$ and $w$.
@@
The number of linear programming constraints and variables Foreplan creates are polynomial in the state space.
@@
By Eq. 2, Foreplan generates one variable per possible state and one constraint for each state and action combination. ∎
Plugging Theorem 5 into Theorem 5 leads to: {theorem} The runtime of Foreplan is polynomial in the number of objects and exponential in $c$ and $w$.
Therefore, with Foreplan, we have introduced an exact planner with a runtime polynomial instead of exponential w.r.t domain sizes for concurrent actions. Thus, we have already achieved an exponential speed up.
Our main contribution, Foreplan, significantly advances the state of the art. However, as it is an exact algorithm, the value function is still not factorized. In case the runtime is of utter importance, we can also approximate the value function, following the idea of approx_factored_mdps, which has also been picked up by, e.g., 10.5555/3020336.3020398. By approximating the value function, we can prevent iterating over the joint state space leading to a blazingly fast, but approximate, version of Foreplan.
@@ 6. Foreplan: Even Faster by Approximation
While Foreplan runs in time polynomial in the number of objects, the runtime still exponentially depends on $c$. In this section, we present an approximation technique inspired by the Approximate Linear Programming (ALP) approach approx_factored_mdps to prevent iterating the whole state space. Our approach follows the same idea as ALP and ALP for first-order MDPs 10.5555/3020336.3020398. We first describe the approximation idea and then how Foreplan uses the approximation for our new representation and for concurrent actions. Last, we give bounds on the runtime and on the approximation quality. We call the approximate version Approximate Foreplan. Since Approximate Foreplan is an extension of our main contribution Foreplan, we refer to the Appendix for technical details.
Foreplan needs to iterate over the whole state space, because the value function maps from a state to the value of that state. We approximate the value function by a set $h_{i}$ of basis functions, whose scope is a subset of $\mathbf{S}$, $V\approx\sum_{i}w_{i}h_{i}$, where the goal is to find the most suitable weights $w_{i}$ approx_factored_mdps. Approximate Foreplan also needs the value of all possible next states in terms of the same approximation. Thus, Approximate Foreplan uses backprojections $g_{i}^{a}$ of the basis functions $h_{i}$ approx_factored_mdps, stating the influence of $x$ on the next state:
@@
To compute the basis functions and backprojections in a lifted way, we parameterize the basis functions analogously to the reward functions and calculate them in the same way. The basis functions should capture the important dynamics in the model koller1999computing, most importantly the rewards 10.5555/3020336.3020398. As proposed by koller1999computing and picked up in 10.5555/3020336.3020398, we also use basis functions for capturing the rewards alongside a constant basis function:
@@
The backprojections are computed in a lifted way:
@@
Approximate Foreplan precomputes all backprojections and then builds the following linear program approx_factored_mdps, whose result is the approximated value function and thus constitutes the result:
@@
The $\alpha_{i}$’s are effectively coefficients for a linear combination over the $w_{i}$, stating how important the minimization of each $w_{i}$ is approx_factored_mdps; doi:10.1287/opre.51.6.850.24925. The maximum operator is no operator in linear programs and is removed in an operation similar to variable elimination (VE) ve. In Appendix A.6, we provide an example for the removal procedure. Moreover, in Appendix B, we show a walkthrough of Approximate Foreplan on the epidemic example.
@@
Most notably, $w$ and induced width in Theorem 6 are mostly small and fixed, leading to a polynomial runtime, as the growth in the number of objects is of more interest. Combining the relational cost graph and the cost networks in a single total relational cost graph, we show in Appendix A.8 that the runtime of Approximate Foreplan is polynomial in the number of objects when the treewidth of the total relational cost graph is bounded.
Moreover, we can show that Approximate Foreplan and ALP compute the same solutions:
{theorem}
@@
The full proof is in Appendix A.9. With appropriate $\alpha_{i}$, the objective function carries out lifted computation of the grounded basis functions. Each constraint is correct, because the lifted backprojections and lifted basis functions compute the same value as for the grounded model. The action representation covers the whole action space. ∎
With ALP and Approximate Foreplan being equivalent, we transfer the approximation guarantee for ALP to Approximate Foreplan: